OpenStudy (anonymous):

Raffle Tickets One thousand raffle tickets are sold for $1 each. One prize of $800 is to be awarded. a) Rena Condos purchases one ticket. Determine her expected value. b) Determine the fair price of a ticket.

OpenStudy (anonymous):

One thousand raffle tickets are sold for $1 each. One prize of $800 is to be awarded. a) Rena Condos purchases one ticket. Find her expected value. The random variable has the values +800 and -1 The corresponding probabilities are 1/1000 and 999/1000 -------------------------------------- Multiply each value by its probability and add the results to get: E(x) = 800(1/1000) + (-1)(999/1000) E(x) = 0.8 - 0.999 E(x) = -0.199 or a loss of 19.9 cents ------------------------------------------ b) Determine the fair price of a ticket. Let the price be "x". The random variable values are 800 and -x The probabilities are 1/1000 and 999/1000 Multiplying and add you get: 800(1/1000)+x(999/1000) = 0 800 + 999x = 0 x = -(800/999) dollars x = -0.80 dollars The price to play should be 80 cents.

OpenStudy (anonymous):

Thank you so very much. This stuff blows my mind.

OpenStudy (anonymous):

yeah, it is a little tricky :P glad I could help